• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name C. Hyde-Wright

    Log entry time 01:43:40 on January16,2002

    Entry number 80353

    keyword=Shift Summary

    1. At start of shift, we still have vacuum problems in the scattering chamber. The experts are looking at it in a controlled access.

    2. 01:20 Dave Meekins reports target must be warmed, extracted and repaired
    All three loops leak. RGA shows H2, He, H2O etc.

    Doug H. concurs.
    night shift standby (except TO until target warm)
    day shift 16-Jan-2002 STANDBY